OpenStudy (anonymous):

Solve the following initial value problem dy/dt + y/(100+t) = 5 y(0) = 0 Hints: y(100) = 1500/4

OpenStudy (anonymous):

Can I use the separable method here?

OpenStudy (anonymous):

use the hint: dy/dt+1500/200*4=5 this means that dy/dt at t=100 is dy/dt=5-15/8 from here you get \(y=\int( 5-15/8 )dt + C\)